OpenStudy (anonymous):

Simplify quantity of negative 3 times x plus 9 all divided by negative 3

jimthompson5910 (jim_thompson5910):

Hint: factor out a negative 3

OpenStudy (ghazi):

\[\frac{ -3x+9 }{ -3 }= -3(x+3)/-3\]